OpenStudy (anonymous):

Limit of x-3/((root of x+6)-3) as x approaches 3?

OpenStudy (freckles):

try multiply the conjugate of the denominator on top and bottom

OpenStudy (anonymous):

would the conjugate be \[\sqrt{x+6}+3\] ?

OpenStudy (freckles):

yep!

OpenStudy (anonymous):

thank you!

OpenStudy (freckles):

so you already got the limit?

OpenStudy (freckles):

\[\lim_{x \rightarrow 3}\frac{x-3}{\sqrt{x+6}-3} \\ \lim_{x \rightarrow 3}\frac{(x-3)(\sqrt{x+6}+3)}{(\sqrt{x+6}-3)( \sqrt{x+6}+3)}\] and you should see: \[(\sqrt{x+6}-3) (\sqrt{x+6}+3)=(x+6)-9=x-3\] so a common factor from top and bottom will "cancel"

OpenStudy (anonymous):

And then the answer would be 6?

OpenStudy (freckles):

sqrt(3+6)+3 sqrt(9)+3 3+3 yes 6

OpenStudy (anonymous):

Thank you!

OpenStudy (freckles):

np

OpenStudy (idku):

\[\lim_{x \rightarrow 3}\dfrac{x-3}{\sqrt{x+6}-3}\]when you plug in x=3, you get 0/0, thus you can differentiate on top and bottom. \[\lim_{x \rightarrow 3}{\Large \frac{1-0}{\frac{1}{2\sqrt{x+6}}-0}}\] \[\lim_{x \rightarrow 3}{\Large \frac{1}{\dfrac{1}{2\sqrt{x+6}}}}\] \[\lim_{x \rightarrow 3}2{\sqrt{x+6}}\]\[2{\sqrt{3+6}}=6\]